OpenStudy (azureilai):

let Q (x,y) and P (3,4) be points on the circle x^2+y^2=25. Find the slope, m, of the line joining P and Q in terms of x.

OpenStudy (anonymous):

x^2+y^2=25 do u know how to find its diff ?

OpenStudy (anonymous):

i mean do u kno how to diffrentiate it ?

OpenStudy (azureilai):

nope. I know how to differentiate linear equations, but not circle equations

OpenStudy (anonymous):

will \(\large x^2+y^2=25\) \(\large y=\sqrt{25-x^2}\)

OpenStudy (azureilai):

oh ok. So then would I rewrite the coordinates for Q as\[(x,\sqrt{25-x^2})\] and solve for the slope as y/x ?

OpenStudy (azureilai):

using point P with Q

OpenStudy (anonymous):

yess

OpenStudy (azureilai):

Would the slope be\[\frac{ 4-\sqrt{25-x^2} }{ 3-x }\]

OpenStudy (anonymous):

but not y/x dy/dx (derivative)

OpenStudy (azureilai):

ok. thank you so much for your help.

OpenStudy (anonymous):

np :)